Mother is four times old as Mary. Five years ago, she was seven times as old. How old will each be in five years?
Answer by mathsolverplus(88)

Mother's age = x
Mary's age = y
Present: x = 4y Mother is four times old as Mary
Past: x-5 = 7(y-5) Both age needs to be subtracted from five first.
4y-5=7y-35
-3y=-30
y=10
x=40 and y=10
Mother's age in five years = 40+5 =45 years old
Mary's age in five years = 10+5 = 15 years old

For more additional instant FREE math help, please check out and follow us on social media!!
Instagram (@mathsolver.plus) : https://www.instagram.com/mathsolver.plus/
Twitter (@MathSolverPlus): https://twitter.com/MathSolverPlus
Snapchat (@mathsolverplus)
Email: mathsolverplus@gmail.com